OpenStudy (anonymous):

simplify: x2+2x+1 ------------------ over 2x+2

OpenStudy (anonymous):

need to factor the top. x^2 + 2x + 1

OpenStudy (anonymous):

x^2 + 2x + 1 = (x + 1)(x+1)

OpenStudy (anonymous):

Now factor out a 2 from the bottom. 2x + 2 = 2(x+1)

OpenStudy (anonymous):